1. Why does water pressure increase with depth?

2. Does pressure depend on density?


1
Expert's answer
2021-02-07T19:10:43-0500

"p=\frac{F}{S}\text{ where}"

"F- \text{weight of the fluid column above the body}"

"F=mg"

"m= V\rho \text{ where }"

"\rho-\text{fluid density}"

"V - \text{liquid column volume, calculated as the volume of a straight cylinder}"

"V=Sh"

"h - \text{cylinder height or otherwise immersion depth}"

"p=\frac{F}{S}=\frac{mg}{S}=\frac{V\rho{g}}{S}=\frac{\rho{gh}}{Sh}=\rho{gh}"

"p=\rho{gh}\ (1)"

"1. \text{from the formula } p=\rho{gh}\text{ follow the increase in pressure with depth}"

"2. \text{from the formula } p=\rho{gh}\text{ pressure is directly proportional to fluid density}"